User Tools

Site Tools


operations:agenda2019_07_11

Chair: Earl

Next week's chair: Jay

Attended: Lim, Jay, Jamie, Ross, Katie, Tiege, Gabor,

Apologies: Simin, Mas, Momtaz,

Absent: Lucas,

General Issues

  • ops8 restarted unexpectedly at 17:45UT and 17:55UT during R4901 (Lim) (foot knocking the power bottom perhaps.)
  • Rostering for next month soon, so please have availabilities listed by the end of this week

Handover Issues

  • CFR111:

Katherine 12m

  Module in Katherine not quite big enough to fit experiment, ~300 GB off. Hopefully there are some missed scans or 'leakage'.
  The sched file late out, maybe?, size was guessed (6.3 rather than 3) Lost an hour or so of experiment.
  New group scheduling, distribution issues.

Yarragadee 12m

A word to all: Yarragadee is currently undergoing testing and changes still, so be extra careful with setups, since the telescope may be left in a different state. Aka DBBC software, FS software (coming soon?) and mark5b/flexbuff.

Tsys still an issue, problem with communcation between boards.

  New autocorrection plots not being uploaded from pcfsyg to ftp server. Most recent plot on vnc screen.
      Update, they now seem to be successfully uploading. Not going to question it (Lucas).
  TSYS giving out complete garbage in BBCs. IF tsys is reasonable and autos (on pcfsyg!) look good (Lucas).
  Every now and then, BBC12 bungs out and the autocorrection fails. It looks like when the plotting system fails completely and you only get solid rainbow colours, but just in that channel (Lucas).
  
  Guifre testing parallel recording. So check for sensible answers when setting up.
  
  TPI overflow (communication problem) related issue. Requires software (dbbc control, ops8 uses version 104_2) reboot. Fails at 5-pt check, has a hard stop (unlike tsys etc.). 
  
  Yg maser VNC IP address indirect. (Brett's working on the networking.(Brett-working)) Currently external to VPN...
  
  Webcam (ptzh...) issues with Yg and Ho... (No zoom?)(Flash updates... web-browser support...)
  
  
  Ke sky-cam may be problematic... or at least, occasionally falls over.
  
  • AUA054

Hobart 26m

  17:30UT Experiment started okay (Momtaz).
  11:30 UT L-band receiver selected change to S/X (Warren) (UT or LT?)
  
  Wrong receiver on axis for first 18 (?) (or 8) hours (74 scans were lost) and wind stows several times during second half.
  LBA experiments use L-band testing? Or Jim P. using it on Monday and didn't change back.
  
  Auto-correlations etc. should still be valid. rxp etc. output live page incorrect.
  
  00:15 UT Record first scan disk_pos/2695871332352,2690890547200,2690922598400 aua054_ho_no0074 (w)
   03:06 UT Wind stow (w)

03:55 UT recorded scans wind stow affected aua054_ho_no0106 through aua054_ho_no0115 (w) 04:00 UT halt schedule, wind here to stay (w) 06:00 UT still wind towed (w) 06:50 UT attem* 06:00 UT still wind stowed (w)pted to restart schedule in lull, wind stowed immediately (w) 07:00 UT something/someone trying to transfer data from module? mk5 reboot, fmset (w) 08:00UT still wind stowed (Lim) 09:09UT schedule resumed around scan 0844-652, wind about 40km/h(Lim) 11:00UT wind speed ~30km/h, still wind stowed. Halt schedule again (Lim) 1150UT wind speed ok, resume schedule (Lim) 1225UT Wind stowed again, halt schedule (Lim) 1347UT Notice that it was still on halt, so continued the schedule (wind was low). Missed scans, no0205 - no0220. (Katie) 1354UT Go the warning: “Cannot change schedules while recording, use_disk_record=off to stop.” (Katie)

Ke

  frequent slewing errors from “onsource” in preob, but tracking soon thereafter. not enough time scheduled between scans-or is antenna slow? (Warren)
  07:40 UT error tc -301 continuous cal not enabled for dbbc ddc solved by changing cont_cal=on in the prc file (Lim)
  
   See above scheduling issue. New scheds overly-optimised, FS update ought to fix.

Yg

   frequent ERROR m5 -906 MARK5 return code 6: inconsistent or conflicting request - noted in recent sessions (w)
   See above scheduling issue.
   
  the timestamp and plot for the auto_correlation spectra are wrong (Lim).

rd1906

Hobart 26m

  1815UT Was getting the 'previous source not reached before new source commanded' error after every scan. Wind was low so no recent wind stows. Noticed that another fs output after every scan was 'Antenna drives not on, switching them on'. Checked the live page and drives were off.Trying using the onsource command but that did not work. Solution was to go to vdesk (type in newsmerd) then 'abort' and then 'drvon'. This clears any previous commands that the antenna gets stuck on. (Katie)
  
  First scan 191-1855 (Katie).
  
  0220 - Ho26 windstowed (JMc)

Live page status showed slewing, drives off before reaching destination. (Drives off but not parked.) Fix via vdesk abort command to cancel last operation then start drives. Error unclear. (Source=disable maybe?)

/home/www/auscope/opswiki/data/pages/operations/agenda2019_07_11.txt · Last modified: 2019/07/11 03:30 by Earl Lester